The director of research and development is testing a new medicine. She wants to know if there is evidence at the 0.05 level that the medicine relieves pain in more than 306 seconds. For a sample of 30 patients, the mean time in which the medicine relieved pain was 316 seconds. Assume the population standard deviation is 19. Find the P-value of the test statistic.
Answer #1

H0: = 306

Ha: > 306

Test statistics

z = ( - ) / ( / sqrt( n ) )

= ( 316 - 306) / ( 19 / sqrt(30) )

= 2.88

p-value = P(Z > z)

= P(Z > 2.88)

= 1 - P(Z < 2.88)

= 1 - 0.9980

= 0.0020